Home‎ > ‎Reports‎ > ‎Crew110a‎ > ‎

day07


Crew #110A Log Book for January 5, 2012

Photos of the day





Commander's Check-In Report

Date and time: 6 January 2012, 20:15

Written by: Julie Mason, Crew 110, Commander


Crew Physical Status: Great


Time Departed/Returned from EVA:  EVA 1 15:00-17:00, Telescope EVA (Last Night): 20:00 – 05:00


Brief Narrative of Field Mission Results:


Last night four crew members imaged Sirius, Capella, Rigel, and Saturn at the Musk Mars Desert Observatory at one hour time increments over the course of the night. The crew processed images using RegiStax 6 today.

Crew members continued to exercise in the habitat while not on EVA, as part of a habitat exercise study.

During EVA 1, three crew members worked on taking images for another 3D image of terrain in the area.


Engineering/Hab Maintenance: Suggest muffling water pump


Report Transmission Schedule:


Plans for Tomorrow:  Welcome Crew 110B


Inventory: Need more oil for the ATV’s and dish soap


Support Requested: None


Miscellaneous: The crew saw people on our EVA who asked to take pictures of us.


Commander's Report

Date and time: 6 January 2012, 20:15
Written by Julie Mason, Crew 110A, Commander

Last night was exciting for the crew astronomers as we imaged Saturn and its magnificent rings with the telescope starting at 3:00am.

During the day, crew members exercised and focused on final reports and one last EVA. A crew photo was taken in the engineering tunnel.

Leaving mars tomorrow will be bittersweet, but I hope that we can come back with new and continued experiments in the future.

Thanks for a great mission.


3D Model Project Report

Date and time: 6 January 2012, 20:21
Written by Aaron Olson, Mission Specialist

During the final day of Crew 110A’s stay at the MDRS site, Will, Julie and I rode ATVs to a few remote areas to gather images to create 3D Models of a few hills that were about 15 minutes away from the habitat. I haven't finished working with the "main" habitat area model, but will continue to work on the model tonight and after our crew leaves Mars on Saturday.

EVA 10 Report

Date:  6 2012
Time: 20:30
Written By: Julie Mason, Crew 110a, Commander

Time of EVA: 15:00-17:00

Crew Members: Aaron Olson, Will Yu, Julie Mason

Site Location: 1/2 mile off of Lowell Highway

Transit Mode: ATV, mostly foot

Objectives: Gather images for another 3D model of surrounding terrain

Results: Collected images for a hill in the area.

Lessons Learned: Smaller hills work best for a single EVA. The images require ability to reach the top of the hill. Preplanning hills to image works best.

Engineering Report

Date and time: 06 January, 2012, 18:00 MST
Written by Will Yu, Crew 110A, Chief Engineer.

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): .75
Propane Tank (Full/0.75/0.5/0.25/Empty): .50
Gasoline Tank (Full/0.75/0.5/0.25/Empty): .50

Oil:
ATV Oil Quantity (# of Quarts in storage): 0
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 12
Trailer Potable Water Tank Level (inches from bottom): 26
Hab Potable Water Tank Level (inches from bottom): 15
Potable Water Meter Reading: 48662.3
Into GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A

Observatory:
Telescope Used? (Yes/No): YES
Observatory Notes/Comments: Sometimes power draw exceeds limit and everything will shut down. Should this happen, reset power surge protector (black power slitter)

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): NO
Oil Checked (yes/no): NO
Tire Status: Good
Odometer at end of day: 125488.9
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: YES
Oil Checked: NO
Fuel Consumed: .3 gal
Tire Status: Good - 4psi
Today's Run Time: 00:30
Season's Run Time: 46:40
Comments:

ATV 2 (Spirit):
Used: YES
Oil Checked: NO
Fuel Consumed: .3 gal
Tire Status: Good - 4psi
Today's Run Time: 00:30
Season's Run Time: 38:03
Comments: Oil level at half full

ATV 3 (Viking I):
Used: YES
Oil Checked: NO
Fuel Consumed: .3 gal
Tire Status: Good
Today's Run Time: 00:30
Season's Run Time: 43:25
Comments: Has broken tail light cover, tail light still works.

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): Upstairs furnace not turned on since new downstairs furnace was enabled
Actual temperature upstairs (Farenheit): 78 F
Thermostat setting downstairs (Farenheit): Downstairs furnace was on during 01:00 – 04:00, 06:00 – 09:30
Actual temperature downstairs (Farenheit): 68 F

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:

General Engineering Notes/Comments:
~ Helmet 5 had an existing crack which has no propagated a little further (7 inches long)
~ Need ATV oil for available for storage
~ ATV3 has broken tail light cover, tail light still works.
~ If back brakes don’t engage, there may be a loose brake cable. First symptom will be difficulty in shifting gears. Re-tighten the cable by the wing nut near the axle and the brakes should work fine.
~ Laptop in the observatory tends to turn off occasionally, which will result in losing the 2 webcam for hab exterior. Simply turn on laptop to restore webcams.

Journalist Report

Date and time: 6 January 2012 7:00 PM

Written by: Mark Ruff, Crew 110a, Mission Specialist/ Journalist

Today was our last day on Mars. It’s been on everyone’s minds as we went about finishing our projects, and making preparations for the next crew, who are likely getting settled for the night in Grand Junction, right now.

We woke at the usual times and made breakfast. We processed data and made lunch. Finally, Julie, Lyndsey, and Will went out for one last EVA, after which we all suited up to take a last set of pictures. Tonight we’ll start packing up our gear, and by noon tomorrow, a team of students from Texas will be here, and we’ll be on our way home to Wisconsin.

The atmosphere in the hab today was a bit strange. Mainly, we were all eager to get back to civilization, even after such a short time away. However, at the same time, this place was our home for a week, and such things are always owed a degree of sentiment.

For most, if not all of us, MDRS was a chance to live out our ambitions of being among the first people to set foot on a new world. For that, we suffered the minor inconveniences of water conservation and dehydrated foods gladly. This is something you don’t get to do every day. There really is nothing like putting on an EVA suit, stepping out the airlock, and thinking that this is what it would feel like to be on Mars. I only hope someday I’ll be able to see for myself.

Chef's Report

Date:6 January 2012

Written by Will Yu, Crew 110A 

1a) Today was pre-prepared day (non-cooking) 

1b) Was today a special day celebrated at the main meal with special food or activity? Today was not a special day, nothing marked a celebration. 

2) List the foods served at today's main meal (usually the evening meal), giving full names of each. Today's main meal was Gourmet Reserves Chicken Rotelle, rehydrated turkey, tortilla soup, and peas 

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded) Some soup leftover 

4) Recipes for anything more complicated than following package directions exactly, or rehydrating. Package directions were followed. Peas and turkey were rehydrated 

5) Lessons learned: Peas take a while to hydrate 

6) Comments/questions for Kim and Jean, the food study investigators 

No comments or questions